OpenStudy (anonymous):

If you do.... A rectangular hole 2cm wide and x cm long is cut in a rectangular sheet of metal 7/2cm wide and y cm long. The length of the hole is 1 cm less than the length of the metal sheet. After the hole is cut, the area of the remaining metal is 11cm squared. Find the length of the hole and the length of the metal sheet. Please let me know, saltreef@gmail.com

OpenStudy (anonymous):

---------Y cm---------- ========== 1 cm _______________________ " " ' 1.5cm " B " ' _"__x cm_____" ' A " " '3.5 cm 2cm " " ' __________"___________" ' 7/2= 3.5 3.5 cm- 2 cm= 1.5 cm

OpenStudy (anonymous):

11 cm sq. - 3.5 cm sq= 7.5 cm sq. Given: Area of Rectangle B is 7.5 cm sq. Length x width (rect. B) = 7.5 cm sq. We know width is 1.5 cm (from 3.5 cm- 2 cm). Solve: Length x width = Area Length x 1.5 = 7.5 Length = 7.5 / 1.5 Length = 5 Therefore: Length of x (hole) = 5 cm. Length of Y (metal sheet) = 5 cm + 1 cm. = 6 c

OpenStudy (anonymous):

kumar, my daughter is in 8th grade. Upon reading your reply, I do not doubt your reply at all, however, I don't know how to explain your reply to her. I don't understand how you came up with the given area as 7.5. If you would, please, please, explain so as my daughter will understand step by step and be able to write it down so she, and I, will understand.